$SPAD/src/input f01mcf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
n:=6
avals:Matrix SF:=
 [[1.0 ,2.0 ,5.0 ,3.0 ,13.0 ,16.0 ,5.0 ,14.0_
 ,18.0 ,8.0 ,55.0 ,24.0 ,17.0 ,77.0 ]]
lal:= 14
nrow:Matrix Integer:= [[1 ,2 ,2 ,1 ,5 ,3 ]]
result:=f01mcf(n,avals,lal,nrow,-1)

———-

References^{[1] nothing}

4